OpenStudy (anonymous):

jonathan and jennifer are testing a physics experiment. jonathan drops a penny from the rooftop of a 160 foot tall building into a pool below on ground level. jennifer, who is next to him, throws her penny straight down at 48 feet per second at the same instant. by how many seconds does her penny beat jonathan's into the water?

OpenStudy (mathstudent55):

\(d = v_0 t + \dfrac{1}{2}gt^2\)

OpenStudy (mathstudent55):

Use the given info each time, and solve for t each time.

OpenStudy (mathstudent55):

For Jonathan, v0 = 0 For Jennifer, v0 = 48 ft/s

OpenStudy (anonymous):

what;s the gravity?

OpenStudy (mathstudent55):

\(d = v_0 t + \dfrac{1}{2}gt^2\) Jonathan: \(-160 = \dfrac{1}{2}(-32.2)t^2\) \(t = 3.15~s\) Jennifer: \(-160 = -48t + \dfrac{1}{2}(-32.2)t^2\) \(16.1t^2 + 48t - 160 = 0\) \(t = \dfrac{-48 \pm \sqrt{48^2 - 4(16.1)(-160)}}{2(16.1)}\) \(t = 1.996\) or \(t = -4.978\) Discard the negative solution, so t = 1.996 for Jennifer.

OpenStudy (mathstudent55):

Now find the difference between the times.